Bob is going to help you out with a series of experiments. You take a blood sample from Bob and find his plasma osmolarity to be 290 mOsm/L. Bob weighs 70 kgs. Bob is going to earn his pay this week. You make him ride an exercise bike in a hot dry room. He loses 500 mOsm electrolytes in 2L of sweat. He also loses 400 mL of water through his respiratory system. Calculate the new plasma osmilarity, new ICF and ECF
Answer #1

Bob is going to help you out with a series of experiments. You take a blood sample from Bob and find his plasma osmolarity to be 290 mOsm/L. Bob weighs 70 kgs. Bob is going to earn his pay this week. You make him ride an exercise bike in a hot dry room. He loses 500 mOsm electrolytes in 2L of sweat. He also loses 400 mL of water through his respiratory system. Calculate the new plasma osmilarity, new ICF and ECF

Answer:-

Sway weighs 70Kg

By and large 60% of the body weight contains water so for 70kg man 60% is roughly 42L water .

The ECF compartment would be 20% o add up to body weight , so it is 14L

The ICF compartment would be 40% of the aggregate body weight , so it is 28L.

Presently new osmolality , ECF and ICF qualities can be determined as pursues.

New osmolality

42L x 290osmoles = 12180 Osm

He lost 2L of water =42 -2 =40L

New osmolality = 12180/40 = 304.5 mosm /L

New ECF

14 Lx 290 osmoles = 4060 mosmoles

New ECF = 4060/304.5 = 13.33 L

New ICF

28L x 290 mOsmoles = 8120 mOsmoles

New ICF = 8120/304.5= 26.71 L

Please do rate answer..........